# Faster cosmological analysis with power spectrum without simulations

### 31 Jan 2024 v2

@@ Abstract
Future surveys could obtain tighter constraints for the cosmological parameters with the galaxy power spectrum than with the Cosmic Microwave Background. However, the inclusion of multiple overlapping tracers, redshift bins, and more non-linear scales means that generating the necessary ensemble of simulations for model-fitting presents a computational burden. In this work, we combine full-shape fitting of galaxy power spectra, analytical covariance matrix estimates, the MOPED compression, and the Taylor expansion interpolation of the power spectrum for the first time to constrain the cosmological parameters directly from a state-of-the-art set of galaxy clustering measurements. We find it takes less than a day to compute the analytical covariance while it takes several months to calculate the simulated ones. Furthermore, the MOPED compression can speed up the posterior sampling by a factor of four. Combined with the Taylor expansion to interpolate the power spectrum, we can constrain the cosmological parameters in just a few hours instead of a few days. Additionally, we find that even without a priori knowledge of the best-fit cosmological or galaxy bias parameters, the analytical covariance matrix with the MOPED compression still gives cosmological constraints consistent, within $0.1\sigma$, after two iterations. The pipeline we have developed here can hence significantly speed up the analysis for future surveys such as DESI and Euclid.
@@ 1 Introduction
In the past few decades, galaxy redshift surveys have become one of the most powerful ways to constrain cosmological parameters. Although their constraining power is still weaker than measurements from the Cosmic Microwave Background (CMB), future galaxy surveys have the potential to surpass the CMB constraints (Carrasco et al., 2012). In a typical analysis pipeline, the redshift measurements from the galaxy surveys are converted to the two-point correlation function or its Fourier counterpart, the power spectrum. Recent developments in the theory of two-point clustering (Ivanov et al., 2020; Tröster et al., 2020; Valogiannis et al., 2020; D'Amico et al., 2021; Chen et al., 2021; Noriega et al., 2022) have allowed us to use galaxy redshift surveys to obtain precise and accurate constraints directly on the parameters of our cosmological model (d'Amico et al., 2020; Glanville et al., 2022; Philcox & Ivanov, 2022; Semenaite et al., 2022), akin to those presented using CMB data.
@@
One caveat of using simulations as described above is that they also introduce noise into the covariance matrix estimation. It is important to note that when using simulations what we actually have is an estimate of the true covariance matrix drawn from a Wishart distribution. As shown in Hartlap et al. (2007); Dodelson & Schneider (2013) and Percival et al. (2014), to suppress the noise in our estimate and generate an unbiased inverse of the covariance matrix, a large number of simulations is required, proportional to the size of the data vector. Furthermore, the appropriate Bayesian approach is to marginalize over our uncertainty of the true covariance matrix which modifies the form of the Gaussian likelihood (Sellentin & Heavens, 2015; Percival et al., 2021). This is a bottleneck for future surveys that will attempt to measure the clustering down to smaller scales and with more redshift bins, and would hence require more simulations. To overcome some of these limitations and generate the covariance matrix faster, semi-analytical (Hamilton et al., 2006; O'Connell et al., 2016; Pearson & Samushia, 2016) and analytical (Sugiyama et al., 2020; Wadekar & Scoccimarro, 2020) covariance matrices have been developed. These methods dramatically reduce the number of simulations required to generate an unbiased and accurate covariance matrix or provide alternative methods to compute the true covariance matrix without simulations.
@@
Alternatively, we can use data compression to reduce the size of the data vector such that we need fewer simulations to estimate its covariance matrix (Heavens et al., 2000; Cannon et al., 2010; Gualdi et al., 2019a). These compression methods can also potentially speed up the data analysis because the compressed covariance matrix and data vector are much smaller. Traditionally, data compression is not expected to significantly speed up the data analysis because the bottleneck is the computation of the power spectrum. However, with recent advancements in machine learning. Several emulators such as Bacco (Arico' et al., 2022) and matryoshka (Donald-McCann et al., 2022) are developed to compute the power spectrum quickly. With these emulators, the bottleneck of the data analysis becomes the chi-squared calculation where data compression would help to speed up the analysis.
@@
In this work, we demonstrate that all of these concepts can be brought together to produce an accurate and fast estimate of the cosmological parameters from galaxy redshift surveys with a minimal number of simulations. We will focus on the analytical covariance matrix from Wadekar & Scoccimarro (2020), the Massively Optimized Parameter Estimation and Data compression (MOPED) algorithm from Heavens et al. (2000), and interpolate the power spectrum with the Taylor expansion method in Colas et al. (2020). All three methods have been applied to power spectrum analysis before (Wadekar et al., 2020; Gualdi et al., 2019b; Colas et al., 2020). However, this is the first time that the combination of all three methods is used to analyze the power spectrum alongside modeling pipelines that allow us to fit directly for cosmological parameters. [^1] In this work, we will use the Python code for Biased tracers in redshift space (PyBird; D'Amico et al. 2021) to calculate the power spectrum using the Effective Theory of Large-Scale Structure; EFTofLSS. A common argument against using these models directly on the two-point clustering is that the large numbers of data points and free parameters used, particularly when fitting multiple redshift bins simultaneously, require a prohibitively large number of simulations for covariance matrix estimation. In this work, we show that by combining the above techniques this is not the case.
@@
This paper is organized as follows. In Section 2, we provide further motivation for this work with simple numerical exercises and a case study involving the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2019). Turning then to real data, Section 3 introduces the combination of redshift data from the six-degree Field Galaxy Survey (6dFGS; Jones et al. 2009), the Baryon Oscillation Spectroscopic Survey (BOSS; Alam et al. 2017) and the extended Baryon Oscillation Spectroscopic Survey (eBOSS; Alam et al. 2021) that we will utilise in this work. In Section 4, we give the theory behind the analytical covariance matrix, the MOPED compression algorithms, and the Taylor expansion method. We present our results in Section 5 before concluding in Section 6.
@@ 2 Motivation
Longer data vectors, or covariance matrices estimated from a small number of simulations, lead to increasingly biased and inaccurate estimations of the likelihood. In the simple case of a multi-variate Gaussian likelihood, Hartlap et al. (2007) and Taylor et al. (2013) developed various analytical formulae to measure the accuracy of the covariance $\operatorname{\boldsymbol{\mathsf{C}}}$ and precision matrices $\operatorname{\boldsymbol{\Psi}}=\operatorname{\boldsymbol{\mathsf{C}}}^{-1}$. For example, the fractional bias can be written (Taylor et al., 2013)
@@
where the unbiased (and unknown) true precision matrix is given by $\hat{\operatorname{\boldsymbol{\Psi}}}$ and the precision matrix as measured from simulations is given by $\operatorname{\boldsymbol{\Psi}}$. $N_{S}$ denotes the number of simulations, and $N_{D}$ is the length of the data vector. [^2] There is a typo in equation (35) of Taylor et al. (2013), the right-hand side of the equation is missing ”-1”. To recover an unbiased estimate of the precision matrix, we hence need to multiply it by the Hartlap factor, $H_{p}$.
@@
$\Delta C_{l_{1},l_{2}}(k_{i},k_{j})$ $=\frac{1}{\sqrt{N_{s}-1}}$ (4)
$\left(C^{2}_{l_{1},l_{2}}(k_{i},k_{j})+C_{l_{1},l_{1}}(k_{i},k_{i})C_{l_{2},l_{2}}(k_{j},k_{j})\right)^{\frac{1}{2}}$
assuming the variations of the power spectra from the mock catalog are Gaussian distributed.
Lastly, Percival et al. (2014) demonstrated that the presence of noise in the covariance matrix should also enlarge parameter constraints obtained using that covariance matrix. They provided an expression for increasing the quoted uncertainty to take this into account if a full Bayesian marginalization over this uncertainty has not been performed (i.e., Sellentin & Heavens 2015). The appropriate factor in this case is
@@
Equations (1)-(5) demonstrate that the fractional bias or error is smaller if the length of the data vector is reduced or the number of simulations is increased. However, they also demonstrate a potential bottleneck in the analysis of next-generation surveys. Take, for instance, the DESI survey configuration presented in DESI Collaboration et al. (2016). This survey aims to produce 3D clustering measurements in $0.1$ width redshift bins between $1.90<z<$ 1.9. In several of these redshift bins, there are three distinct, overlapping tracers, meaning one might aim to simultaneously fit a joint data vector containing separate, but correlated, auto-clustering measurements for each of these tracers. Furthermore, due to differing observational systematics in the Northern and Southern Galactic Caps, previous surveys such as BOSS (Beutler et al., 2016) and eBOSS (Alam et al., 2021) have often treated the clustering in these two sky-patches as separate, but sometimes correlated, data vectors. Combined, this means we could expect to fit a total data vector consisting of up to 6 correlated auto-clustering statistics in each redshift bin (for now we will ignore the fact we may also want to measure and fit the cross-clustering between these tracers or sky-patches).
@@
Based again on the same previous surveys, we can assume each set of clustering statistics could contain $\sim 50$ measurement bins (monopole and quadrupole of the power spectrum between $0\,h\,\mathrm{Mpc}^{-1}<k<0.20\,h\,\mathrm{Mpc}^{-1}$, plus hexadecapole between $0\,h\,\mathrm{Mpc}^{-1}<k<0.10\,h\,\mathrm{Mpc}^{-1}$, all with $0.01\,h\,\mathrm{Mpc}^{-1}k$-bins). The length of the joint data vector for a single redshift bin could hence be on the order of 300 correlated measurements. Taking a standard EFTofLSS-style approach (d'Amico et al., 2020; Glanville et al., 2022), we would fit this data vector with a model containing $\sim 6$ cosmological parameters, plus $10$ bias parameters per Galactic Cap, per tracer (so $66$ free parameters in total). The punchline is that to avoid having to increase our parameter uncertainties by more than $10\%$ ($5\%$) due to covariance matrix noise, we would need $\sim 2000$ (3600$\sim$ 3600)∼ 3600 ) simulations for such a redshift bin. Although not impossible, this is clearly computationally demanding and may need to be repeated for other redshift bins or as a result of other systematic tests/validations within the collaboration.
@@
Figure 1: This plot shows the fractional error (bottom) and fractional bias (top) for the precision matrix for data from the BOSS galaxy survey as a function of the number of simulations used in the estimation. In both cases, red points/lines indicate the numerical/analytical results for the full precision matrix, while blue points/lines represent the numerical/analytical results for the compressed precision matrix. Both panels show a clear improvement when using compression.
@@
One way to reduce this burden is to use the MOPED algorithm. We demonstrate this using power spectrum covariance matrices from the BOSS survey (Beutler & McDonald 2021; detailed more in Section 3) estimated using up to 2048 simulations. We estimate the standard deviation and mean of the precision matrix from these simulations using the jackknife method. The numerical result is shown in Fig 1, in excellent agreement with the predictions from equations (1) and (2). These results also show that the compressed precision matrix has a factor of 10 (2) improvement in the fractional bias (error) compared to its uncompressed counterpart for the same number of simulations, motivating its further exploration in this work.
@@ 3 Data
For tests in this work, we used the combined datasets of the 6-degree Field Galaxy Survey (6dFGS; Jones et al. 2009), the $z1$ and $z2$ redshift bins of the Baryon Oscillation Spectroscopic Survey Data Release 12 (BOSS DR12; Alam et al. 2017) for both NGC and SGC sky-patches, and the LRGpCMASS and QSO samples of the extended Baryon Oscillation Spectroscopic Survey Data Release 16 (eBOSS DR16; Alam et al. 2021) for both NGC and SGC sky-patches. The combined sample of these 9 correlated clustering sample creates one of the largest consistent datasets for testing cosmological models; their respective mock catalogs are introduced in Beutler & McDonald (2021) and the dataset itself is presented thoroughly in section 5 therein and in section 3.1 of Glanville et al. (2022) and so not repeated here. For all of these data, we account for the window function by convolving the model with a matrix form of the window function, computed as in Beutler & McDonald (2021).
We only analyze the monopole and quadrupole of the power spectrum. The range of wavevectors we have chosen to fit is $0.01h\mathrm{Mpc}^{-1}\leq k\leq 0.20h\mathrm{Mpc}^{-1}$ for all data sets. The result is a single joint data vector consisting of 342 measurement bins. As a ballpark number, from our fiducial fitting method we find that approximately half of our cosmological information comes from the BOSS data.
@@ 4.1 Analytical covariance matrix
The analytical covariance matrix (Wadekar & Scoccimarro, 2020) is developed with the SPT approach while PyBird is an EFT model. Although it is possible to map bias parameters between these two bases, the PyBird model invokes some extra assumptions such around the “UV subtraction" (Perko et al., 2016) compared to other EFT approaches. Consequently, the number of bias parameters in PyBird is less than that in the analytical covariance matrix. Therefore, direct mapping is not possible. However, we know “$b_{1}$" in SPT and EFT both denote the galaxy bias, we can use the local Lagrangian relations to estimate other bias parameters.
@@
$C^{\mathrm{Anal}}$ $=C_{G}(P_{l})+C_{\mathrm{BC}}(b_{1}^{*},b_{2}^{*},\gamma_{2}^{*})+C_{\mathrm{LA}}(b_{1}^{*},b_{2}^{*},\gamma_{2}^{*})+$ (7)
@@
$P_{l}$ $=\sum b_{G,i}P_{l,\mathrm{lin},i}(b_{1},b_{2},b_{4})+$ (8)
$P_{l,\mathrm{const}}(b_{3},c_{ct},c_{r,1},c_{r,2},c_{\epsilon,1},c_{\epsilon,\mathrm{mono}},c_{\epsilon,\mathrm{quad}}).$
Here, we use *** to denote the bias parameters in the SPT basis and the ones without *** are in the EFT basis. We also only highlight the dependence of the analytical covariance matrix ($C^{\mathrm{Anal}}$) and the nonlinear model of power spectrum from PyBird ($P_{l}$) on the bias parameters since they are the main focus in this section. They also depend on the cosmological parameters and other properties such as number density and FKP weight. The full expression for the Gaussian covariance ($C_{G}$), Local-Averaging covariance ($C_{\mathrm{LA}}$), Beat-Coupling covariance ($C_{\mathrm{BC}}$), regular trispectrum ($C_{T_{0}}$), and shot-noise ($C_{\mathrm{SN}}$) are given in equation (57), (77), (62), (66), and (91, 92) of Wadekar & Scoccimarro (2020) respectively. The expression for $P_{l,\mathrm{lin},i}(b_{1},b_{2},b_{4})$ and $P_{l,\mathrm{const}}(b_{3},,c_{ct},c_{r,1},c_{r,2},c{\epsilon,0},c{\epsilon,1},c{\epsilon,2})$ are given in equation (30) of d'Amico et al. (2020).
Equation (4.1) demonstrates we need to know $b^{u}={b_{1}^{*},b_{2}^{*},\gamma_{2}^{*},b_{3}^{*},\gamma_{3}^{*},\gamma_{2}^{x,*},\gamma_{21}^{*}}$ to calculate the analytical covariance matrix. Since we know $b_{1}^{*}=b_{1}$, we can use the local Lagrangian relation (Eggemeier et al., 2019) to express the non-local SPT bias parameters in terms of $b_{1}$:
@@
For the local bias parameters in the SPT, we find them with the empirical fitting formulae in Lazeyras et al. (2016). Their final expressions are
@@
Comparing to the equations in Lazeyras et al. (2016), equation (13) has an extra $-\frac{8}{21}(b_{1}-1)$ and equation (14) has an extra $\frac{796}{1323}(b_{1}-1)-\frac{8}{7}b_{2}^{*}$. This is because Lazeyras et al. (2016) is using a different bias basis than Wadekar & Scoccimarro (2020), these extra factors convert the local bias parameters into the correct bias basis. Now, all parameters needed to calculate the analytical covariance matrix can be written as a function of $b_{1}$. We then substitute them into the formulae in Wadekar & Scoccimarro (2020) to find the analytical covariance matrix.

### 05 Apr 2024 v3

@@ Abstract
Future surveys could obtain tighter constraints on the cosmological parameters with the galaxy power spectrum than with the Cosmic Microwave Background. However, the inclusion of multiple overlapping tracers, redshift bins, and more non-linear scales means that generating the necessary ensemble of simulations for model-fitting presents a computational burden. In this work, we combine full-shape fitting of galaxy power spectra, analytical covariance matrix estimates, the MOPED (Massively Optimised Parameter Estimation and Data compression) method, and the Taylor expansion interpolation of the power spectrum for the first time to constrain the cosmological parameters directly from a state-of-the-art set of galaxy clustering measurements. We find it takes less than a day to compute the analytical covariance while it takes several months to calculate the simulated ones. Combining MOPED with the Taylor expansion interpolation of the power spectrum, we can constrain the cosmological parameters in just a few hours instead of a few days. We also find that even without a priori knowledge of the best-fit cosmological or galaxy bias parameters, the analytical covariance matrix with the MOPED compression still gives consistent cosmological constraints to within $0.1\sigma$ after two iterations. Therefore, the pipeline we have developed here can significantly speed up the analysis for future surveys such as DESI and Euclid.
@@ 1 Introduction
In the past few decades, galaxy redshift surveys have become one of the most powerful ways to constrain cosmological parameters. Although their constraining power is still weaker than measurements from the Cosmic Microwave Background (CMB), future galaxy surveys have the potential to surpass the CMB constraints (Carrasco et al., 2012). In a typical analysis pipeline, the redshift measurements from the galaxy surveys are converted to the two-point correlation function or its Fourier counterpart, the power spectrum. Recent developments in the theory of two-point clustering (McEwen et al., 2016; Simonović et al., 2018; Ivanov et al., 2020; Tröster et al., 2020; Valogiannis et al., 2020; D'Amico et al., 2021; Chen et al., 2021; Noriega et al., 2022) have allowed us to use galaxy redshift surveys to obtain precise and accurate constraints directly on the parameters of our cosmological model (D'Amico et al., 2020; Glanville et al., 2022; Philcox & Ivanov, 2022; Semenaite et al., 2022), akin to those presented using CMB data.
@@
One caveat of using simulations as described above is that they also introduce noise into the covariance matrix estimation. It is important to note that when using simulations, what we actually have is an estimate of the true covariance matrix drawn from a Wishart distribution. As shown in Hartlap et al. (2007); Dodelson & Schneider (2013) and Percival et al. (2014), to suppress the noise in our estimate and generate an unbiased inverse of the covariance matrix, a large number of simulations is required, proportional to the size of the data vector. Furthermore, the appropriate Bayesian approach is to marginalize over the uncertainty of the true covariance matrix, which modifies the form of the Gaussian likelihood (Sellentin & Heavens, 2015; Percival et al., 2021). This is a bottleneck for future surveys that will attempt to measure the clustering down to smaller scales and with more redshift bins, so we would require more simulations. To overcome some of these limitations and generate the covariance matrix faster, semi-analytical (Hamilton et al., 2006; O'Connell et al., 2016; Pearson & Samushia, 2016), approximations (Monaco, 2016; Blot et al., 2019), and analytical (Sugiyama et al., 2020; Wadekar & Scoccimarro, 2020) covariance matrices have been developed. These methods dramatically reduce the number of simulations required to generate an unbiased and accurate covariance matrix or provide alternative methods to compute the true covariance matrix without simulations.
@@
Alternatively, we can use data compression to reduce the size of the data vector such that we need fewer simulations to estimate its covariance matrix (Heavens et al., 2000; Cannon et al., 2010; Gualdi et al., 2019a). These compression methods can also potentially speed up the data analysis because the compressed covariance matrix and data vector are much smaller. Furthermore, with recent advancements in machine learning, several emulators such as Bacco (Pellejero Ibañez et al., 2023) and matryoshka (Donald-McCann et al., 2022) significantly speed up the power spectrum computation. The data analysis bottleneck becomes the chi-squared computation when using the emulators. In this case, the data compression can help to speed up the analysis because the size of the compressed data vectors and covariance matrix are smaller.
@@
In this work, we demonstrate that all of these concepts can be brought together to produce an accurate and fast estimate of cosmological parameters from galaxy redshift surveys with minimum numbers of simulations. We will focus on the analytical covariance matrix from Wadekar & Scoccimarro (2020), the Massively Optimized Parameter Estimation and Data compression (MOPED) algorithm from Heavens et al. (2000), and the Taylor expansion interpolation of the power spectrum in Colas et al. (2020). All three methods have been applied to power spectrum analysis before (Wadekar et al., 2020; Gualdi et al., 2019b; Colas et al., 2020). However, this is the first time that the combination of all three methods is used to analyze the power spectrum alongside modeling pipelines that allow us to fit directly for cosmological parameters. In this work, we will use the Python code for Biased tracers in redshift space (PyBird; D'Amico et al. 2021) to calculate the power spectrum using the Effective Theory of Large-Scale Structure; EFTofLSS. A common argument against using Full-Shape modelling such as PyBird directly on the two-point clustering is that the large numbers of data points and free parameters used, particularly when fitting multiple redshift bins simultaneously, require a prohibitively large number of simulations for covariance matrix estimation. We can show that this is not the case by combining the above techniques in this work.
@@
This paper is organized as follows. In Section 2, we provide further motivation for this work with simple numerical exercises and a case study involving the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2019). Turning then to real data, Section 3 introduces the combination of redshift data from the six-degree Field Galaxy Survey (6dFGS; Jones et al. 2009), the Baryon Oscillation Spectroscopic Survey (BOSS; Alam et al. 2017) and the extended Baryon Oscillation Spectroscopic Survey (eBOSS; Alam et al. 2021) that we will utilize in this work. In Section 4, we introduce the theory behind the analytical covariance matrix, the MOPED compression algorithms, and the Taylor expansion interpolation of the power spectrum. We present our results in Section 5 before concluding in Section 6.
@@ 2 Motivation
Longer data vectors, or covariance matrices estimated from a small number of simulations, lead to increasingly biased and inaccurate likelihood estimation. In the simple case of a multi-variate Gaussian likelihood, Hartlap et al. (2007) and Taylor et al. (2013) developed various analytical formulae to measure the accuracy of the covariance $\operatorname{\boldsymbol{\mathsf{C}}}$ and precision matrices $\operatorname{\boldsymbol{\Psi}}=\operatorname{\boldsymbol{\mathsf{C}}}^{-1}$. For example, the fractional bias can be written (Taylor et al., 2013) as
@@
where the unbiased (and unknown) true precision matrix is given by $\hat{\operatorname{\boldsymbol{\Psi}}}$ and the precision matrix as measured from simulations is given by $\operatorname{\boldsymbol{\Psi}}$. $N_{S}$ denotes the number of simulations, and $N_{D}$ is the length of the data vector. [^1] There is a typo in equation (35) of Taylor et al. (2013), the right-hand side of the equation is missing ”-1”. Therefore, to recover an unbiased estimate of the precision matrix, we need to multiply it by the Hartlap factor, $H_{p}$.
@@
$\Delta C_{l_{1},l_{2}}(k_{i},k_{j})$ $=\frac{1}{\sqrt{N_{s}-1}}$
$\left(C^{2}_{l_{1},l_{2}}(k_{i},k_{j})+C_{l_{1},l_{1}}(k_{i},k_{i})C_{l_{2},l_{2}}(k_{j},k_{j})\right)^{\frac{1}{2}}$ (4)
assuming the variations of the power spectra from the mock catalog are Gaussian distributed. Here, $l_{1}$ and $l_{2}$ denote the multipole of the covariance matrix.
Lastly, Percival et al. (2014) demonstrated that the noise in the covariance matrix should also enlarge parameter constraints obtained using that covariance matrix. They provided an expression for increasing the quoted uncertainty to take this into account if a full Bayesian marginalization over this uncertainty has not been performed (i.e., Sellentin & Heavens 2015). The appropriate factor in this case is
@@
Equations (1)-(5) demonstrate that the fractional bias or error is smaller if the length of the data vector is reduced or the number of simulations is increased. However, they also demonstrate a potential bottleneck when analyzing next-generation surveys. Take, for instance, the DESI survey configuration presented in DESI Collaboration et al. (2016). This survey aims to produce 3D clustering measurements in $0.1$ width redshift bins between $1.90<z<$ 1.9. In several of these redshift bins, there are three distinct overlapping tracers. It means one might aim to simultaneously fit a joint data vector containing separate but correlated, auto-clustering measurements for each tracer. Furthermore, due to differing observational systematics in the Northern and Southern Galactic Caps, previous surveys such as BOSS (Beutler et al., 2016) and eBOSS (Alam et al., 2021) have often treated the clustering in these two sky-patches as separate, but sometimes correlated, data vectors. Combined, this means we could expect to fit a total data vector consisting of up to 6 correlated auto-clustering statistics in each redshift bin (we will ignore the fact we may also want to measure and fit the cross-clustering between these tracers or sky-patches for now).
@@
Based again on the same previous surveys, we can assume each set of clustering statistics could contain $\sim 50$ measurement bins (monopole and quadrupole of the power spectrum between $0\,h\,\mathrm{Mpc}^{-1}<k<0.20\,h\,\mathrm{Mpc}^{-1}$, plus hexadecapole between $0\,h\,\mathrm{Mpc}^{-1}<k<0.10\,h\,\mathrm{Mpc}^{-1}$, all with $0.01\,h\,\mathrm{Mpc}^{-1}k$-bins). Therefore, the length of the joint data vector for a single redshift bin could be on the order of 300 correlated measurements. Taking a standard EFTofLSS-style approach (D'Amico et al., 2020; Glanville et al., 2022), we would fit this data vector with a model containing $\sim 6$ cosmological parameters, plus $10$ bias parameters per Galactic Cap, per tracer (so $66$ free parameters in total). The punchline is that to avoid having to increase our parameter uncertainties by more than $10\%$ ($5\%$) due to covariance matrix noise, we would need $\sim 2000$ (3600$\sim$ 3600)∼ 3600 ) simulations for such a redshift bin. Although not impossible, this is clearly computationally demanding and may need to be repeated for other redshift bins or as a result of other systematic tests/validations within the collaboration.
@@
Figure 1: This plot shows the fractional error (bottom) and fractional bias (top) for the precision matrix with data from the BOSS galaxy survey as a function of the number of simulations used in the estimation. In both cases, red points/lines indicate the numerical/analytical results for the full precision matrix, while blue points/lines represent the numerical/analytical results for the compressed precision matrix. Both panels show a clear improvement when using compression.
@@
One way to reduce this burden is to use the MOPED algorithm. We demonstrate this using power spectrum covariance matrices from the BOSS survey (Beutler & McDonald 2021; detailed more in Section 3) estimated using up to 2048 simulations. We estimate the standard deviation and mean of the precision matrix from these simulations with the jackknife method. The numerical result shown in Fig 1 is in excellent agreement with the predictions from equations (1) and (2). These results also show that the compressed precision matrix has a factor of 10 (2) improvement in the fractional bias (error) compared to its uncompressed counterpart for the same number of simulations, motivating its further exploration in this work.
@@ 3 Data
For tests in this work, we used the combined datasets of the 6-degree Field Galaxy Survey (6dFGS; Jones et al. 2009), the $z1$ and $z2$ redshift bins of the Baryon Oscillation Spectroscopic Survey Data Release 12 (BOSS DR12; Alam et al. 2017) with both NGC and SGC sky-patches, and the LRGpCMASS and QSO (Quasi Stellar Object) samples of the extended Baryon Oscillation Spectroscopic Survey Data Release 16 (eBOSS DR16; Alam et al. 2021) with both NGC (North Galactic Cap) and SGC (South Galactic Cap) sky-patches. The combination of these nine correlated clustering samples creates one of the largest consistent datasets for testing cosmological models; their respective mock catalogs are introduced in Beutler & McDonald (2021) and the dataset itself is presented thoroughly in section 5 therein and section 3.1 of Glanville et al. (2022) and so not repeated here. For all of these data, we account for the window function by convolving the model with a matrix form of the window function, computed as in Beutler & McDonald (2021).
We only analyze the monopole and quadrupole of the power spectrum. The range of wavevectors we have chosen to fit is $0.01h\mathrm{Mpc}^{-1}\leq k\leq 0.20h\mathrm{Mpc}^{-1}$ for all data sets. The result is a single joint data vector consisting of 342 measurement bins. As a ballpark number, from our fiducial fitting method, we find that approximately half of our cosmological information comes from the BOSS data.
@@ 4.1 Analytical covariance matrix
The analytical covariance matrix (Wadekar & Scoccimarro, 2020) uses a different bias basis than the one in PyBird. Although it is possible to map bias parameters between these two bases, the PyBird model invokes some extra assumptions such as the “UV subtraction" (Perko et al., 2016) compared to other EFT approaches. Consequently, the number of bias parameters in PyBird is less than that in the analytical covariance matrix. Therefore, direct mapping is not possible. However, we know “$b_{1}$" in SPT and EFT both denote the linear galaxy bias, so we can use the local Lagrangian relations to estimate other bias parameters. For simplicity, we will denote the bias parameters in the analytical covariance matrix as the “SPT bias parameters" and the bias parameters in PyBird as “EFT bias parameters".
@@
$C^{\mathrm{Anal}}$ $=C_{G}(P_{l})+C_{\mathrm{BC}}(b_{1}^{*},b_{2}^{*},\gamma_{2}^{*})+C_{\mathrm{LA}}(b_{1}^{*},b_{2}^{*},\gamma_{2}^{*})+$
@@
$P_{l}=\sum b_{G,i}P_{l,\mathrm{lin},i}(\vec{b}_{\mathrm{NG}})+P_{l,\mathrm{const}}(\vec{b}_{G})$ (8)
and
$\vec{b}_{G}$ =$=\{b_{3},c_{ct},c_{r,1},c_{r,2},c_{\epsilon,1},c_{\epsilon,\mathrm{mono}},c_{\epsilon,\mathrm{quad}}\},$
$\vec{b}_{\mathrm{NG}}$ =$=\{b_{1},b_{2},b_{4}\},$
$\vec{b}$ $=\vec{b}_{G}\cup\vec{b}_{\mathrm{NG}}.$ (9)
Here, we use *** to denote the bias parameters in the bias basis of the analytical covariance matrix, and the ones without *** are in the bias basis of PyBird. We also only highlight the dependence of the analytical covariance matrix ($C^{\mathrm{Anal}}$) and the nonlinear model of power spectrum from PyBird ($P_{l}$) on the bias parameters since they are the main focus in this section. They also depend on the cosmological parameters and other properties such as number density and FKP weight. The full expression for the Gaussian covariance ($C_{G}$), Local-Averaging covariance ($C_{\mathrm{LA}}$), Beat-Coupling covariance ($C_{\mathrm{BC}}$), regular trispectrum ($C_{T_{0}}$), and shot-noise ($C_{\mathrm{SN}}$) are given in equation (57), (77), (62), (66), and (91, 92) of Wadekar & Scoccimarro (2020) respectively. The expression for $P_{l,\mathrm{lin},i}(\vec{b}_{G})$ and $P_{l,\mathrm{const}}(\vec{b}_{\mathrm{NG}})$ are
$P_{l,\mathrm{const}}=P_{l}(k,\vec{b})|_{\vec{b}_{G}\rightarrow 0}$ (10)
and
$P_{l,\mathrm{lin},i}=\frac{\partial P_{l}(k,\vec{b})}{\partial b_{G,i}}|_{\vec{b}_{G}\rightarrow 0}.$ (11)
Equation (7) demonstrates we need to know $b^{u}=\{b_{1}^{*},b_{2}^{*},\gamma_{2}^{*},b_{3}^{*},\gamma_{3}^{*},\gamma_{2}^{x,*},\gamma_{21}^{*}\}$ to calculate the analytical covariance matrix. Since we know $b_{1}^{*}=b_{1}$, we can use the local Lagrangian relation (Eggemeier et al., 2019) to express the non-local SPT bias parameters in terms of $b_{1}$:
@@
For the local SPT bias parameters, we find them with the empirical fitting formulae in Lazeyras et al. (2016). Their final expressions are
@@
Comparing to the equations in Lazeyras et al. (2016), equation (16) has an extra $-\frac{8}{21}(b_{1}-1)$ and equation (17) has an extra $\frac{796}{1323}(b_{1}-1)-\frac{8}{7}b_{2}^{*}$. This is because Lazeyras et al. (2016) uses a different bias basis than Wadekar & Scoccimarro (2020), and these extra factors convert the local bias parameters into the correct bias basis. Now, all parameters needed to calculate the analytical covariance matrix can be written as a function of $b_{1}$. We then substitute them into the formulae in Wadekar & Scoccimarro (2020) to find the analytical covariance matrix.